OpenStudy (mejia941):

A coffee mixture has beans that sell for $0.72 a pound and beans that sell for $0.52. If 110 pounds of beans create a mixture worth $0.74 a pound, how much of each bean is used?

OpenStudy (benji):

Okay, let's let the $0.72 a pound mixture be represented as x, and let's let the $0.52 be represented with y. We know that the two mixtures have to add up to 110 lbs at $0.74 a pound so \[110*0.74=81.4\] With this information we can write the equation\[0.72x+0.52y=81.4\] So now we know \[x+y=180\] or \[x=-y+180\]And we can plug this x into the first equation to get\[0.72(-y+180)+0.52y=81.4\]What do you think the next step is?

OpenStudy (mejia941):

im confused

OpenStudy (benji):

What are you confused on?

OpenStudy (mejia941):

where did you get the 180 from???

OpenStudy (benji):

I meant 110 not 180, it's going to be the total pounds we have.

OpenStudy (benji):

So with the equation \[0.72(−y+110)+0.52y=81.4\]What's the next step?

OpenStudy (mejia941):

well i did the equation and got 79=81.4

OpenStudy (abdiweli):

How did you get 79 please